OpenStudy (anonymous):

Evaluate the function f(x)= 4 • 7^x for x = -1 and x = 2. Show your work. Now is it asking to be down in two different ways like f(x)= 4 • 7^-1 and f(x)= 4 • 7^2 If so, did I do them right... f(x)= 4 • 7^-1 4 • 1/7^1 4/7 f(x)= 4 • 7^2 4 • 49 196 As always any help would be grateful. Thanks.

OpenStudy (danjs):

you have it right

OpenStudy (danjs):

what are you asking?

OpenStudy (anonymous):

When I first saw the question I thought it was asking to put a number in the (x) spot but I went back to look in my book and never saw it done that way. Just wanted to ask to be safe to know it was done the right way and not to make a fool of myself.

OpenStudy (campbell_st):

well another way it to know a little about powers when multiplying add the powers \[f(2) = 4\times 7^2 = 4 \times 7^{-1} \times 7^3= \frac{4}{7} \times 7^3\] which yields the same answer

OpenStudy (campbell_st):

you method seems fine, based on the information

OpenStudy (anonymous):

Ok,Good. Thanks for the help.